OpenStudy (anonymous):

log base 5 (x+1) = 3 solve

OpenStudy (anonymous):

https://www.youtube.com/user/Hairfunfashion4teens

OpenStudy (anonymous):

\[Log_ab=c--->a^c=b\]

OpenStudy (anonymous):

\[alogb=\log(b^a)\]

OpenStudy (anonymous):

If the base is not expressed, it is 10.

OpenStudy (anonymous):

5 (x+1) = 3 5x+5=3 -5 -5 5x=-2 5x/5=-2/5 x=-2/5

OpenStudy (anonymous):

i think that right

OpenStudy (anonymous):

Sammy that's wrong, it is logarithms.

OpenStudy (anonymous):

okay sorry

OpenStudy (anonymous):

\[Log_5(x+1)=3\]

OpenStudy (anonymous):

I know the answer is 124 because i got it right but idk how to show the work to get the answer

OpenStudy (anonymous):

\[5^{3}=x+1\]\[125=x+1\]\[x=124\]YEP

OpenStudy (anonymous):

Thank you i will have many more coming

OpenStudy (anonymous):

You Welcome.